OpenStudy (chris215):

The surface area, S, of a sphere of radius r feet is S = S(r) = 4πr2. Find the instantaneous rate of change of the surface area with respect to the radius r at r = 4.

OpenStudy (superdavesuper):

instantaneous rate of change of the surface area S with respect to the radius r = dS/dr = d(4πr2)/dr = ...

OpenStudy (chris215):

whats d stand for?

OpenStudy (superdavesuper):

d stands for differentiation...this is calculus right?

OpenStudy (chris215):

yeah

OpenStudy (superdavesuper):

so u should know about how to differentiate a function like 4πr2...

OpenStudy (chris215):

i got 16pi

OpenStudy (superdavesuper):

dats not the right ans...wuld u show ur steps plz?

OpenStudy (chris215):

4π(4^2)=64pi/4=16pi

OpenStudy (superdavesuper):

dats not the right steps...sorry! plz follow the ones below: instantaneous rate of change of the surface area S with respect to the radius r = dS/dr = d(4πr^2)/dr = 4π * d(r^2)/dr = ?

OpenStudy (radar):

Don't forget that after you differentiate, substitute the value 4 for "r"

OpenStudy (radar):

Please show what you got for the dS/dr (the derivative)

OpenStudy (radar):

Hint: Use the "power rule"

OpenStudy (radar):

I am gonna have to run. Good luck getting the derivative.

OpenStudy (chris215):

8πr

OpenStudy (superdavesuper):

GREAT @chris215 yes, dS/dr = 8πr now all u need to do is to put r=4 to find the rate of change at that point :)

OpenStudy (chris215):

oh ok so 32pi

OpenStudy (superdavesuper):

Correct @chris215 good job! :)

OpenStudy (chris215):

thanks so much